Sunland Bunyon Lumber Co. produces several products that can be sold at the split-off point or processed further and then sold. The following results are from a recent period:

Product Sales Value
at Split-off
Additional
Variable Costs
Sales Value after
Further Processing
Green lumber $152600 $23200 $175800
Rough lumber 117000 26700 163800
Sawdust 97000 18500 123600


Which products should be processed further?

All three products.

Rough lumber and sawdust.

Green lumber and rough lumber.

Green lumber and sawdust.

**Hello, I am not sure on how the formula works on this. Thank you!
Answer #1

this question based on sell or process further .

ROUGH LUMBER:

sale value at split off +additional variable cost= 117000+26700

=143700

sale value after further processing =163800

GREEN LUMBER:

sale value at split off +additional variable cost =152600+23200

=175800

sale value after further procesing is =175800

SAWDUST:

sale value at split off +additional variable cost =97000+18500

=15500

sale value after further processing is =123600

HERE IN THIS PROBLEM WE HAVE TO COMPARE SALE VALUE + ADDITIONAL VARIABLE COST WITH SALE VALUE AFTER FURTHER PROCESSING .AND WE HAVE TO SELECT WHICH IS GIVE MORE PROFIT .

A) ROUGH LUMBER AND SAWDUST;

rough lumber profit =163800-143700

=20100

sawdust profit =123600-115500

=8100

*here rough lumber gives more profit when compared to sawdust .so,process further rough lumber.

B) GREEN LUMBER AND ROUGH LUMBER;

green lumber profit = 175800-175800

=0

rough lumber profit = 163800-143700

=20100

*here rough lumber gives more profit when compared to green lumber.so, process further rough lumber.

C) GREEN LUMBER AND SAWDUST;

green lumber profit =175800-175800

=0

sawdust profit =123600-115500

=8100

* here sawdust give more profit when compared to green lumber .so, process further sawdust.